OpenStudy (anonymous):

find the slope and y intercept of the line 14x+4y=24

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (cp9454):

slope = \[\frac{ -coefficient of x }{ coefficient of y }\]

OpenStudy (cp9454):

slope = \[\frac{ -14 }{ 4 }\]

OpenStudy (anonymous):

slope= -7/2 and the y intercept is 6